OpenStudy (anonymous):

1.What is the length of rectangle with an area of 48 yd2 and a width of 3 yd?

OpenStudy (welshfella):

area = width * length 45 = 3 * L ( where L = length) can you find the value of L?

OpenStudy (anonymous):

Is it 16

OpenStudy (welshfella):

no divide both sides by 3 so L = 45/3

OpenStudy (welshfella):

oh yes sorry - i wrote 45 instead of 48 its 48/3 = 16 yes